Name a point on a given line, then find the slope of the line.
y+2=3/2(x+7)
A. Point:(2,7) Slope: -3/2
B. Point:(7,2) Slope: 3/2
C. Point (-7,-2) Slope: 3/2
D. Point:(-7,-2) Slope: -3/2

Answer by Alan3354(69443)

Name a point on a given line, then find the slope of the line.
y+2=3/2(x+7)
A. Point:(2,7) Slope: -3/2
B. Point:(7,2) Slope: 3/2
C. Point (-7,-2) Slope: 3/2
D. Point:(-7,-2) Slope: -3/2
-----------
y+2=3/2(x+7) is the form y-y1 = m*(x-x1)
where m = slope and (x1,y1) is a point.
===========================
The long way:
Step 1, find the slope.
That eliminates 2.
Step 2, sub the values and check.